OpenStudy (anonymous):

3x+2y=3,3x-2y=-9..solve

OpenStudy (anonymous):

so y is 12?

OpenStudy (anonymous):

keep going im learning,

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

okay, got it, keep going please

OpenStudy (nincompoop):

-4y = -12 solve for y

OpenStudy (anonymous):

what did you use anyway

OpenStudy (nincompoop):

divide both sides by -4

OpenStudy (anonymous):

1.5?

OpenStudy (anonymous):

no, i got 3 right

OpenStudy (anonymous):

well 1.5 and 3

OpenStudy (anonymous):

are you still there

OpenStudy (nincompoop):

y = 3 then plug that value into the one of the original equations to solve for x

OpenStudy (anonymous):

so then 4 divided into 12 =3

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

no reverse that 3 divided into 12 x=4

OpenStudy (nincompoop):

3x+2y=3 and y =3 3x + 2(3) = 3 3x + 6 = 3 3x + 6 - 6 = 3 - 6 3x = -3 divide both sides by 3

OpenStudy (nincompoop):

3x - 2y=-9 and y = 3 3x - 2(3) = -9 3x - 6 = -9 3x = -3 divide both sides by 3 you get the same value for x

OpenStudy (anonymous):

what wold we have

OpenStudy (nincompoop):

laughing out loud I gave you the entire solution already

OpenStudy (nirmalnema):

3x+2y=3,3x-2y=-9..